OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. The function notation is f(x)=-2/3x+400. What would the graph for the function notation represent?

OpenStudy (anonymous):

this is simple

OpenStudy (anonymous):

@triciaal Please help! I will be back on about 30 mins.

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\bf 2x+3y=1200\qquad y=f(x)\qquad thus \\ \quad \\ 2x+3f(x)=1200\implies 3f(x)=-2x+1200\implies f(x)=\cfrac{-2x+1200}{3} \\ \quad \\ f(x)={\color{brown}{-\cfrac{2}{3}}}x+400\impliedby {\color{brown}{ slope\implies rate\ of\ change}}\) keep in mind that, that's the rate of change of "y" as "x" fluctuates so if say x = 0 then -2/3 x = 0 and -2/3x + 400 = 0 + 400 that means, that if you produce NO smartphone cases, x = 0, then you'd be able to produce 400 tablet cases and show a profit of 1,200 the more smartphone cases you'd produce, would decrease the tablet cases you can produce, at that rate staying within the 1,200 constaint

OpenStudy (anonymous):

@awesomeness_dude I still dont get what the answer to my question is

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@morganvaughn the question is: The function notation is f(x)=-2/3x+400. What would the graph for the function notation represent? y = number of tablet cases x = number of phone cases f() = y so the equation is like saying \( y = -\frac{2}{3}x + 400 \) x is the number of phones cases so if we input 0 for x that means y = 400 and we can sell 400 tables cases and still make a profit of 1,200. If x = 3 that means we sold three phone cases and y would equal 398, which means if we sell 3 phone cases we should not sell / produce more than 398 tables cases to stay at within 1,200 profit.

OpenStudy (anonymous):

so @MDoodler is "x is the number of phones cases so if we input 0 for x that means y = 400 and we can sell 400 tables cases and still make a profit of 1,200. If x = 3 that means we sold three phone cases and y would equal 398, which means if we sell 3 phone cases we should not sell / produce more than 398 tables cases to stay at within 1,200 profit." the answer or you explaining?

OpenStudy (anonymous):

I get it. Thank you!